# KSHSeek: Data-Driven Approaches to Mitigating and Detecting Knowledge-Shortcut Hallucinations in Generative Models

### v1

@@ Abstract
Zhiwei Wang1, Zhongxin Liu2, Ying Li1, Hongyu Sun3, Meng Xu4, Yuqing Zhang1,2,3 1University of Chinese Academy of Sciences, 2Xidian University, 3Hainan University, 4University of Waterloo wangzhiwei23@mails.ucas.ac.cn
@@ 1 Introduction
The emergence of large language models (LLMs) has brought a paradigm shift to natural language processing (NLP), especially in generative tasks such as question-answering (Rangapur and Rangapur, 2024; Michail et al., 2023; Qin et al., 2023) However, this revolution has also caused a growing concern, known as model hallucinations. Huang et al.Huang et al. (2024) building on the definition of hallucinations proposed by Ji et al.(Dziri et al., 2021; Ji et al., 2023), expanded the applicability and scope of the term, classifying model hallucinations into two types: factual hallucinations and faithfulness hallucinations. This expanded classification provides a new paradigm for understanding model hallucinations.
@@
A major cause of the aforementioned hallucinations is knowledge-shortcut(Ju et al., 2024;Li et al., 2022). The training data often contains a significant amount of information such as computer system version numbers and book indexes. LLMs have learned the comparative features of this data and erroneously applied these features to the comparison of regular numbers, leading to hallucinations. In the classification proposed by LiLi et al. (2022), this cause is referred to as a "Knowledge Shortcut". Building on this concept, we define hallucinations caused by knowledge shortcuts as knowledge-shortcut hallucinations.
Knowledge shortcut arises because language models typically do not genuinely understand the intricate and complex factual knowledge but rather rely on shortcuts. They tend to over-rely on semantically proximate positions in the pre-training data, shared high-frequency words, and the quantity of related documents Kandpal et al. (2023).This can introduce spurious correlation biases, causing the model to produce hallucinations even when working with correct and defect-free data sources.
@@
A common approach for mitigating data-related hallucinations is data filtering, including strictly controlling data source(Gao et al., 2020; Gunasekar et al. (2023)) and deduplication. Deduplication which is divided into exact and near duplicates faces challenges. Exact duplicate detection is inefficient for large datasets Manber and Myers (1993), while near duplicate method like hash-based algorithm MinHash Broder (1997) prioritize speed but miss hidden information. Semantic duplicate recognition using pre-trained models Abbas et al. (2023) is slower and impractical for large datasets. Thus, balancing granularity in duplicate detection and processing speed, while effectively reducing knowledge-shortcut hallucinations, remains a challenge.
This paper focuses on analyzing knowledge-shortcut hallucinations triggered by high-similarity texts from correct, defect-free data. We propose a High Similarity Pruning Algorithm that mitigates knowledge-shortcut hallucinations from a data perspective by leveraging semantic similarity, shared high-frequency words. Furthermore, based on these characteristics and incorporating model uncertainty(Xiao and Wang, 2021; Miao et al., 2023), we design a hybrid detection method tailored for CQA tasks to identify knowledge-shortcut hallucinations effectively. Our mitigation strategy demonstrates promising results across multiple LLMs and parameter scales. Notably, in the fine-tuning of nanoGPT-large, it successfully reduces knowledge-shortcut hallucinations by 6.5%.
@@ 2.2 Data High Similarity Pruning
1) For each row in $data_{1}$, compute the top $K_{1}$ Jaccard and TF-IDF similarity values with the remaining $(n-1)$ datasets. Record the corresponding indices and values. 2) Calculate the top $K_{2}$ most frequent indices (High-Frequency group, $G_{HF}$) and the top $K_{2}$ largest values (High-Value group, $G_{HV}$). 3) Combine the four groups (HF and HV for both Jaccard and TF-IDF similarities), remove duplicates, and identify rows in $data_{j\in n}$ to delete, denoted as $R_{1,j\in n}$(Equation 1):
@@ 2.3 Detection of Knowledge-Shortcut Hallucination
Detecting knowledge-shortcut hallucinations requires distinguishing them from other hallucinations. We focus on fact-based question-answering tasks with a CQA structure, where the correct answer is embedded in the context. To detect knowledge-shortcut hallucinations, we propose a method combining similarity features and self-check uncertainty measurementMiao et al. (2023). The pseudocode is in Appendix B.3.
@@ 3.1.1 Datasets
We selected four datasets with a CQA structure from the generative text datasets on Hugging Face as training or fine-tuning datasets $(data_{1},data_{2},data_{3},data_{4})$, along with a hallucination test dataset as the ablation test dataset. The four CQA datasets belong to different domains, forming a diverse training or fine-tuning datasets. In terms of data quantity selection, not all data from the four datasets were used. The large discrepancies in the total volume of data across different datasets could make it difficult for the model to learn long-tail knowledgeKandpal et al. (2023), thus negatively impacting the experimental results. Therefore, we selected portions of data from trivia-cqa [^2] https://huggingface.co/datasets/tilyupo/trivia_cqa and QASports [^3] https://huggingface.co/datasets/PedroCJardim/QASports that closely matches the sample sizes of sciq [^4] https://huggingface.co/datasets/allenai/sciq and financial-qa-10K [^5] https://huggingface.co/datasets/virattt/financial-qa-10K. Details of this data selection can be found in Table 2.
@@ 3.1.2 Model Selection
We conducted our experiments on two generative models which is distributed under the MIT License. We used the model according to the terms specified in the license: nanoGPT [^7] https://github.com/karpathy/nanoGPT and TinyLlama [^8] https://github.com/jzhang38/TinyLlama. For nanoGPT, we selected three parameter scales: gpt2-large (774M), gpt2-medium (350M), and gpt2 (124M), to perform fine-tuning and training experiments. For TinyLlamaZhang et al. (2024), we conducted fine-tuning experiments using the LoRAHu et al. (2021) (Low-Rank Adaptation) method at the 1.1B parameter scale.
@@ 4 Related Works
Significant progress has been made in the study of hallucinations in LLMs. XuXu et al. (2024) argue that eliminating hallucinations in LLMs is impossible. Numerous techniques have emerged to mitigate LLM hallucinations, with notable approaches including Retrieval-Augmented Generation (RAG)Lewis et al. (2020), knowledge retrievalVarshney et al. (2023), CoNLILei et al. (2023), and CoVeDhuliawala et al. (2023). Our aim is to clarify the underlying causes of knowledge-shortcut hallucinations and minimize such hallucinations as much as possible.
Shortcut learning is a critical area of research in LLM hallucinations, with knowledge shortcuts representing the manifestation of shortcut learning at the data level. GeirhosGeirhos et al. (2020) and DuDu et al. (2023) suggest that dataset bias is the starting point of shortcut learning, and many excellent works have focused on alleviating shortcut learning from the data perspective, such as identifying biased sentencesLeiLei et al. (2022), data shortcutsFriedman et al. (2022), or replacing datasets with more balanced onesTang et al. (2023). Tang’s researchTang et al. (2021) indicates that fine-tuned language models can learn and even amplify biases present in the training datasets, leading to poor performance in downstream tasks, which aligns with our experimental findings. Despite these advancements, existing works have yet to fully explore all the ways in which dataset bias can manifest. Our goal is to conduct an in-depth study of one type of hallucination triggered by correct, defect-free data sources in large models and propose a feasible method for mitigating and predicting such illusions.
@@ B.3 Pseudocode of Mitigation and Detection
8:end for$(data_{1},...data_{n})^{\prime}\leftarrow(data_{1},data_{n})-R_{all}$;

### v2

@@ Abstract
Zhongxin Liu1, Zhiwei Wang2, Jun Niu1, Ying Li2, Hongyu Sun3, Meng Xu4, He Wang1, Gaofei Wu1, Yuqing Zhang1,2,3 1Xidian University, 2University of Chinese Academy of Sciences, 3Hainan University, 4University of Waterloo liuzhongxin@stu.xidian.edu.cn
@@ 1 Introduction
The emergence of large language models (LLMs) has brought a paradigm shift to natural language processing (NLP), especially in generative tasks such as question-answering (rangapur2024battle; michail2023uzh_clyp; qin2023chatgpt) However, this revolution has also caused a growing concern, known as model hallucinations. Huang et al.Huang:23 building on the definition of hallucinations proposed by Ji et al.(NPH-2021-neural; ji2023survey), expanded the applicability and scope of the term, classifying model hallucinations into two types: factual hallucinations and faithfulness hallucinations. This expanded classification provides a new paradigm for understanding model hallucinations.
@@
A major cause of the aforementioned hallucinations is knowledge-shortcut(ju2024investigating;li2022pre). The training data often contains a significant amount of information such as computer system version numbers and book indexes. LLMs have learned the comparative features of this data and erroneously applied these features to the comparison of regular numbers, leading to hallucinations. In the classification proposed by Lili2022pre, this cause is referred to as a "Knowledge Shortcut". Building on this concept, we define hallucinations caused by knowledge shortcuts as knowledge-shortcut hallucinations.
Knowledge shortcut arises because language models typically do not genuinely understand the intricate and complex factual knowledge but rather rely on shortcuts. They tend to over-rely on semantically proximate positions in the pre-training data, shared high-frequency words, and the quantity of related documents kandpal23.This can introduce spurious correlation biases, causing the model to produce hallucinations even when working with correct and defect-free data sources.
@@
A common approach for mitigating data-related hallucinations is data filtering, including strictly controlling data source(gao2020pile; gunasekar2023textbooks) and deduplication. Deduplication which is divided into exact and near duplicates faces challenges. Exact duplicate detection is inefficient for large datasets manber1993suffix, while near duplicate method like hash-based algorithm MinHash broder1997minhash prioritize speed but miss hidden information. Semantic duplicate recognition using pre-trained models abbas2023semdedup is slower and impractical for large datasets. Thus, balancing granularity in duplicate detection and processing speed, while effectively reducing knowledge-shortcut hallucinations, remains a challenge.
This paper focuses on analyzing knowledge-shortcut hallucinations triggered by high-similarity texts from correct, defect-free data. We propose a High Similarity Pruning Algorithm that mitigates knowledge-shortcut hallucinations from a data perspective by leveraging semantic similarity, shared high-frequency words. Furthermore, based on these characteristics and incorporating model uncertainty(xiao2021hallucination; miao2023selfcheck), we design a hybrid detection method tailored for CQA tasks to identify knowledge-shortcut hallucinations effectively. Our mitigation strategy demonstrates promising results across multiple LLMs and parameter scales. Notably, in the fine-tuning of nanoGPT-large, it successfully reduces knowledge-shortcut hallucinations by 6.5%.
@@ 2.2 Data High Similarity Pruning
1) For each row in $data_{1}$, compute the top $K_{1}$ Jaccard and TF-IDF similarity values with the remaining (n1)(n-1) datasets. Record the corresponding indices and values. 2) Calculate the top $K_{2}$ most frequent indices (High-Frequency group, $G_{HF}$) and the top $K_{2}$ largest values (High-Value group, $G_{HV}$). 3) Combine the four groups (HF and HV for both Jaccard and TF-IDF similarities), remove duplicates, and identify rows in $data_{j\in n}$ to delete, denoted as $R_{1,j\in n}$(Equation 1):
@@ 2.3 Detection of Knowledge-Shortcut Hallucination
Detecting knowledge-shortcut hallucinations requires distinguishing them from other hallucinations. We focus on fact-based question-answering tasks with a CQA structure, where the correct answer is embedded in the context. To detect knowledge-shortcut hallucinations, we propose a method combining similarity features and self-check uncertainty measurementmiao2023selfcheck. The pseudocode is in Appendix B.3.
@@ 3.1.1 Datasets
We selected four datasets with a CQA structure from the generative text datasets on Hugging Face as training or fine-tuning datasets $(data_{1},data_{2},data_{3},data_{4})$, along with a hallucination test dataset as the ablation test dataset. The four CQA datasets belong to different domains, forming a diverse training or fine-tuning datasets. In terms of data quantity selection, not all data from the four datasets were used. The large discrepancies in the total volume of data across different datasets could make it difficult for the model to learn long-tail knowledgekandpal23, thus negatively impacting the experimental results. Therefore, we selected portions of data from trivia-cqa [^2] https://huggingface.co/datasets/tilyupo/trivia_cqa and QASports [^3] https://huggingface.co/datasets/PedroCJardim/QASports that closely matches the sample sizes of sciq [^4] https://huggingface.co/datasets/allenai/sciq and financial-qa-10K [^5] https://huggingface.co/datasets/virattt/financial-qa-10K. Details of this data selection can be found in Table 2.
@@ 3.1.2 Model Selection
We conducted our experiments on two generative models which is distributed under the MIT License. We used the model according to the terms specified in the license: nanoGPT [^7] https://github.com/karpathy/nanoGPT and TinyLlama [^8] https://github.com/jzhang38/TinyLlama. For nanoGPT, we selected three parameter scales: gpt2-large (774M), gpt2-medium (350M), and gpt2 (124M), to perform fine-tuning and training experiments. For TinyLlamazhang2024tinyllama, we conducted fine-tuning experiments using the LoRAhulora (Low-Rank Adaptation) method at the 1.1B parameter scale.
@@ 4 Related Works
Significant progress has been made in the study of hallucinations in LLMs. Xuxu2024hallucination argue that eliminating hallucinations in LLMs is impossible. Numerous techniques have emerged to mitigate LLM hallucinations, with notable approaches including Retrieval-Augmented Generation (RAG)lewis2020retrieval, knowledge retrievalvarshney2023stitch, CoNLIlei2023chain, and CoVedhuliawala2023chain. Our aim is to clarify the underlying causes of knowledge-shortcut hallucinations and minimize such hallucinations as much as possible.
Shortcut learning is a critical area of research in LLM hallucinations, with knowledge shortcuts representing the manifestation of shortcut learning at the data level. Geirhosgeirhos2020shortcut and Dudu2023shortcut suggest that dataset bias is the starting point of shortcut learning, and many excellent works have focused on alleviating shortcut learning from the data perspective, such as identifying biased sentencesLeilei2022sentence, data shortcutsfriedman2022finding, or replacing datasets with more balanced onestang2023large. Tang’s researchtang2021mitigating indicates that fine-tuned language models can learn and even amplify biases present in the training datasets, leading to poor performance in downstream tasks, which aligns with our experimental findings. Despite these advancements, existing works have yet to fully explore all the ways in which dataset bias can manifest. Our goal is to conduct an in-depth study of one type of hallucination triggered by correct, defect-free data sources in large models and propose a feasible method for mitigating and predicting such illusions.
@@ B.3 Pseudocode of Mitigation and Detection
[…]